Each square in the grid is 0.50 m x 0.50 m in dimensions. Force F = 2025 N and a force H =1000N. Find the moments of each force about the points A,B,C,and D.


1
Expert's answer
2021-08-10T08:01:04-0400



"AF=\sqrt 5,"

"BF=\frac 1{\sqrt5},"

"CF=\frac 9{\sqrt5},"

"DF=\sqrt5,"

"AH=\frac6{\sqrt{10}},"

"BH=\frac{12}{\sqrt{10}},"

"CH=\frac8{\sqrt{10}},"

"DH=\frac{11}{\sqrt{10}},"


"M_{AF}=2264~\text N\cdot \text m,"

"M_{BF}=453~\text N\cdot \text m,"

"M_{CF}=4075~\text N\cdot \text m,"

"M_{DF}=2264~\text N\cdot \text m,"

"M_{AH}=949~\text N\cdot \text m,"

"M_{BH}=1897~\text N\cdot \text m,"

"M_{CH}=1265~\text N\cdot \text m,"

"M_{DH}=1739~\text N\cdot \text m."